OpenStudy (anonymous):

Is my answer correct? I'm suppose to write an equivalent expression using radical notation. Problem: 16^3/4 is the answer 2?

jimthompson5910 (jim_thompson5910):

\[16^{\frac{3}{4}}\] \[(\sqrt[4]{16})^3\] \[(2)^3\] \[8\] So \[16^{\frac{3}{4}}=8\] You're close, you just forgot to cube the 2.

OpenStudy (anonymous):

oops. Thanks again! =)